Adjusted WACC.  Lewis runs an outdoor adventure company and wants to know what effect a tax change will have on his​ company's WACC. ​ Currently, Lewis has the following financing​ pattern:

Equity: 42% and cost of 18.66%

Preferred stock: 18% and cost of 12.72%

Debt: 40% and cost of 11.3% before taxes.

What is the adjusted WACC for Lewis is the tax rate is

a) 35%? b) 30%? c) 15%? d) 10%? e) 0%?
Answer #1

WACC=(Weight of equity)*(Cost of equity)+(Weight of Preferred stock)*(Cost of Preferred stock)+(Weight of Debt)*(Cost of Debt before tax)*(1-Tax rate)
Substituting the values except the tax rate, we get;

WACC=(42%)*(18.66%)+(18%)*(12.72%)+(40%)*(11.3%)*(1-Tax rate)
WACC=0.078372+0.022896+0.0452*(1-Tax rate)
WACC=0.101268+0.0452*(1-Tax rate)

Now, we need to determine the WACC for different tax rates

Part a: At a tax rate of 35%
WACC=0.101268+0.0452*(1-35%)
=0.101268+0.02938=0.130648 or 13.06% (Rounded to two decimal places)

Part b: At a tax rate of 30%
WACC=0.101268+0.0452*(1-30%)=0.132908 or 13.29% (Rounded to two decimal places)

Part c: At a tax rate of 15%
WACC=0.101268+0.0452*(1-15%)=0.139688 or 13.97% (Rounded to two decimal places)

Part d: At a tax rate of 10%
WACC=0.101268+0.0452*(1-10%)=0.141948 or 14.19% (Rounded to two decimal places)

Part e: At a tax rate of 0%
WACC=0.101268+0.0452*(1-0%)=0.146468 or 14.65% (Rounded to two decimal places)